# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 7.2 Student’s t‑Regression with $\ell^{0}$-Penalty
All the numerical experiment results of the algorithms for solving (7.21) are illustrated in Table 3. Across all tested scenarios, GCNM consistently exhibits the most efficient performance in terms of the CPU time. It reaches the stopping criterion $\|x^{k}-\widehat{x}^{k}\|\leq 10^{-4}$ significantly faster than both GSSN and ZeroFPR regardless of either the problem size, or the regularization parameter $\mu$. Furthermore, GCNM obtains the smallest values of the residual $\|Ax^{k}-b\|$, often in the order of $10^{-7}$ or lower indicating highly accurate approximate solutions to the equation $Ax=b$. ZeroFPR, while showing slightly larger values of $\|Ax^{k}-b\|$ than GCNM, still maintains reasonably accurate approximations and requires a significantly fewer CPU time than GSSN. It performs especially well at moderate problem sizes and smaller values of $\mu$.
@@ 7.3 Restoration of Noisy Blurred Images
As a nonconvex example, we consider the restoration of a $256\times 256$ cameraman test image. The degradation process involves applying a Gaussian blur with a kernel of size $9\times 9$ and standard deviation $4$, followed by the addition of standard Gaussian noise with standard deviation $10^{-3}$. The restoration is formulated as a nonconvex optimization problem (7.1). We evaluate the performance under different settings of the smooth regularization parameter $\mu_{2}\in\{5\times 10^{-2},5\times 10^{-3}\}$ and of the nonconvex regularization parameter $\mu_{0}\in\{10^{-4},10^{-5}\}$. All methods are initialized at $x^{0}=b$. We evaluate and compare the performance of our proposed GCNM algorithm (Algorithm 4) against two efficient methods for image restoration: the proximal gradient method (PGM), also known as the forward-backward splitting method, as developed in [3, Section 5.1], and the ZeroFPR method with BFGS updates introduced in [58]. We run GCNM, PGM, and ZeroFPR until the stopping criterion $\eta_{k}:=\|x^{k}-\hat{x}^{k}\|\leq 10^{-2}$ is satisfied. The original and blurred images are shown in Figure 1. As illustrated in Figure 2, all three solvers produce visually satisfactory restorations. Table 4 presents the residuals $\eta_{k}$, and the computational time (in seconds) required to reach the prescribed accuracy. Among the three methods, GCNM and ZeroFPR consistently outperform PGM, achieving the target accuracy in smaller time. GCNM yields the smallest residual $\eta_{k}$ within a comparable or even significantly shorter runtime, highlighting its superior efficiency in these scenarios. Figure 3 shows convergence rates of the algorithms in one of the generated problems.
@@
Figure 3: Residual and CPU times of three solvers with $\mu_{0}=10^{-4}$, $\mu_{2}=5\times 10^{-3}$
@@ Acknowledgements
Research of Boris S. Mordukhovich was partly supported by the US National Science Foundation under grant DMS-2204519 and by the Australian Research Council under Discovery Project DP250101112. The authors declare that the presented results are new, and there is no any conflict of interest.

### v3

@@ 7.2 Student’s t‑Regression with $\ell^{0}$-Penalty
All the numerical results for the algorithms applied to (7.18) are reported in Table 3. In the table, “time” denotes the computational time in seconds, “iter” denotes the number of iterations required to reach the prescribed accuracy, $\delta_{k}:=\|Ax^{k}-b\|$, $\eta_{k}:=\|x^{k}-\widehat{x}^{k}\|$, and “nnz” denotes the number of nonzero components in the obtained approximate solution. Across all tested scenarios, GCNM consistently exhibits the most efficient performance in terms of the time. It reaches the stopping criterion $\eta_{k}\leq 10^{-4}$ significantly faster than both GSSN and ZeroFPR regardless of either the problem size, or the regularization parameter $\mu$. Furthermore, GCNM obtains the smallest values of the residual $\delta_{k}$, often in the order of $10^{-7}$ or lower, indicating therefore highly accurate approximate solutions to the equation $Ax=b$. ZeroFPR, while showing slightly larger values of $\delta_{k}$ than GCNM, still maintains reasonably accurate approximations and requires a significantly fewer time than GSSN. It performs especially well at moderate problem sizes and smaller values of $\mu$. From the table, we can observe that the solutions obtained by the different methods differ, as expected for a nonconvex problem. The line search strategy used in GSSN generally requires more iterations, whereas GCNM in these experiments appears to take mostly full Newton steps, resulting in fewer iterations. We present below the following special 2-dimensional example of (7.18), which demonstrates that the methods can converge to different points, potentially minimizers or $M$-stationary points.
Example 7.6.
(illustration of different algorithm performances). Consider the optimization problem (7.18) with the initial data:
$m=1,\;n=2,\;A=[1\quad 1],\;b=1,\;\nu=1,\;\mu=\frac{1}{10},$
which corresponds to the explicit formulation
$\mbox{\rm minimize}\quad\varphi(x_{1},x_{2})=\log\left[1+(x_{1}+x_{2}-1)^{2}\right]+\frac{1}{10}\|(x_{1},x_{2})\|_{0}\quad\text{\rm subject to }\;(x_{1},x_{2})\in{\rm I\!R}^{2}.$ (7.22)
Using the subdifferential formula for the $\ell_{0}$-norm from Proposition 7.1 together with the gradient of the smooth term $f$ in (7.20), we can find the M-stationary points of (7.22) as follows:
$\mathcal{X}_{M}:=\{(x_{1},x_{2})\in{\rm I\!R}^{2}\;|\;(0,0)\in\partial\varphi(x_{1},x_{2})\}=\{(x_{1},x_{2})\in{\rm I\!R}^{2}|\;x_{1}+x_{2}=1\}.$
It is clear that $(1,0)$ and $(0,1)$ are global minimizers of (7.22). We now conduct numerical experiments using our GCNM Algorithm 4 together with other methods including GSSN and ZeroFPR. All the methods are tested from two different starting points and are terminated once the condition $\|x^{k}-\widehat{x}^{k}\|\leq 10^{-4}$ is satisfied. The results are summarized in Table 4. As indicated by the numerical results, both GCNM and GSSN converge to the global optimal solution $(0,1)$ when initialized at $(5,5)$, whereas ZeroFPR converges to the M-stationary point $(1/2,1/2)$. When all algorithms are initialized at (5,5)(-5,5), they converge to the M-stationary point (9/2, 11/2)(-9/2,\,11/2). These observations motivate further investigation into the conditions under which our algorithm converges to a global optimal solution rather than to a merely M-stationary point, a behavior that is guaranteed only under the theoretical assumptions developed in the previous sections.
Table 4: Numerical results for different starting points
starting point method iterations approximate solution
$(5,5)$ GCNM 71 $(0,0.99999977)$
GSSN 18 $(0,0.99999939)$
ZeroFPR 30 $(0.50008385,0.50008385)$
(5,5)(-5,5) GCNM 7 (4.50144135,5.50144135)(-4.50144135,5.50144135)
GSSN 4 (4.51120306,5.51120444)(-4.51120306,5.51120444)
ZeroFPR 21 (4.49992114,5.50007886)(-4.49992114,5.50007886)
@@ 7.3 Restoration of Noisy Blurred Images
We consider the restoration of a $256\times 256$ cameraman test image. The degradation process involves applying a Gaussian blur with a kernel of size $9\times 9$ and standard deviation $4$, followed by the addition of standard Gaussian noise with standard deviation $10^{-3}$. The restoration is formulated as a nonconvex optimization problem (7.1). We evaluate the performance under different settings of the smooth regularization parameter $\mu_{2}\in\{5\times 10^{-2},5\times 10^{-3}\}$ and of the nonconvex regularization parameter $\mu_{0}\in\{10^{-4},10^{-5}\}$. All methods are initialized at $x^{0}=b$. We evaluate and compare the performance of our proposed GCNM algorithm (Algorithm 4) against two efficient methods for image restoration: the proximal gradient method (PGM), also known as the forward-backward splitting method, as developed in [3, Section 5.1], and the ZeroFPR method with BFGS updates introduced in [57]. We run GCNM, PGM, and ZeroFPR until the stopping criterion $\eta_{k}:=\|x^{k}-\widehat{x}^{k}\|\leq 10^{-2}$ is satisfied. The original and blurred images are shown in Figure 1. As illustrated in Figure 2, all three solvers produce visually satisfactory restorations. Table 5 presents the residuals $\eta_{k}$, and the computational time (in seconds) required to reach the prescribed accuracy. Among the three methods, GCNM and ZeroFPR consistently outperform PGM, achieving the target accuracy in smaller time. GCNM yields the smallest residual $\eta_{k}$ within a comparable or even significantly shorter runtime, highlighting its superior efficiency in these scenarios. Figure 3 shows convergence rates of the algorithms in one of the generated problems.
@@
Figure 3: Residual and times of three solvers with $\mu_{0}=10^{-4}$, $\mu_{2}=5\times 10^{-3}$
@@ Acknowledgements
Research of Pham Duy Khanh is funded by Ho Chi Minh City University of Education Foundation for Science and Technology under grant number CS.2025.19.04TD. Research of Boris S. Mordukhovich was partly supported by the US National Science Foundation under grant DMS-2204519 and by the Australian Research Council under Discovery Project DP250101112. Research of Vo Thanh Phat was partly supported by the Early Career Scholars Program 2026, University of North Dakota under project 43700-2375-UND0031286. The authors declare that the presented results are new, and there is no any conflict of interest.